Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
boldrini, capocasa, menzione, nardecchia, rocchi - 22:33 Monday 16 May 2022 (55863) Print this report
PR RH power increase @ 11 W

This afternoon the power applied to the PR RH has been increased up to 11 W at 15.32 UTC in the CITF configuration. 

From the simulations we expect Delta RoC= -7.9 m (RoC=1477 m→ 1469.1 m) and a thermal lens of 5.2e-5 dioptres.

The temperature of the coils have been monitored during the transient to make sure it did not reach 40 degrees C (see fig.1) .

At around 19 UTC, ISC people started to try to lock the ITF @Carm null.

Images attached to this report
Comments to this report:
derossi, spinicelli, chiummo - 23:03 Thursday 19 May 2022 (55915) Print this report

Triggered by the current miscentering of the carrier beam on EPRB B4 PC Cam, we noticed that its position changed when the RH was switched on (plot 1).

Figures 1 and 2 compare the image when the arms are locked on IR (as you can see the carrier is now on the top right). To be checked at carm null.

Images attached to this comment
Search Help